# Predicting Mergers and Acquisitions: Temporal Dynamic Industry Networks

### v2

@@ Abstract
Merger and Acquisition (M&A) has increasing importance over market consolidation and reconstruction. For companies who decide to acquire other firms, M&A is a major investment practice, while one of the major incentives to trigger a M&A is obtaining complementarity to enhance its market power among a competitive local industry. Besides the intrinsic factors that have been widely considered in the existing empirical studies, the M&A behavior of a firm is actively influenced by the M&A behaviors of their peer firms, called “peer effect”. However, existing research fails to capture such rich inter-dependencies of M&A events in the industry network. Also, a favorable M&A predictive model should be able to make fine-grained deal-level predictions and do not require ad-hoc feature engineering. This kind of model can provide precise predictions of rival firms’ M&A behaviors and M&A recommendations for specific bidder companies and target companies. Yet most existing works can only predict the behavior of one side of MA deals, which fail to make deal recommendations specific to a firm. Plus, most existing predictive works also require ad-hoc truncations of the data, which transfer continuous time information to ad-hoc truncations. The information loss caused by transformation is harmful to predictive performance. Because the transformation is laborious and arbitrary, this makes those models hard to apply in the real-world scenario. Finally, due to the sparsity of MA events, existing deal-level predictive models require ad-hoc data rebalance which brings bias to the model. This may further harm the performance and reliability of the predictions. To address these limitations, we propose a MA predictive model based on Temporal Dynamic Industry Network (TDIN). Our method is based on temporal point processes and deep learning methods. The deep learning models enable our model to capture the rich inter-dependencies among MA events. Our method is also able to make fine-grained deal-level predictions without ad-hoc feature engineering and data rebalance: the temporal point processes model the sparsity of events inside its intensity function. Evaluation results show the superiority of our method, on the basis of the MA cases between January 1997 to December 2020.
@@ 2.2 Temporal Point Processes
We have the relationship between the intensity function and probability density function333proof see Appendix:
@@ 2.3 Graph Neural Network
The receptive field can be increased by stacking aggregation and combination mechanisms. If we have 2 layers of them, for every node, it can aggregate the information over 2-hop neighbors. However, generally speaking, the number of layers cannot be larger than 2 444see section 4.4 of (Hamilton et al., 2017) for detail. Aggregate and combine components are also called ”message passing mechanism”. Formally, consider the graph $G=(V,E)$, where $V$ is the set of nodes. The feature matrix of nodes is $X$. The aggregate and combine components at layer $l$ can be represented by:
@@ 3.2.1 Timing Module
Specifically, for the vector representation of intrinsic factors $e_{d}^{t}$, we consider two parts of the information as the input: accounting variables and textual embeddings from 10-K555only the “product description” part of 10-K. The reasons why we consider textual information are: first, recent studies have shown that textual information from 10-K has significant predictive power for M&A events (Hoberg & Philips, 2010; Hoberg & Philips, 2016; Routledge et al., 2017). Secondly, one of the most important motivations of firms continuing to conduct M&A transactions is that they are seeking product synergy to maintain competitive advantage (Hoberg & Philips, 2010). However, accounting variables have no capability to provide adequate information about firms’ products. In the meanwhile, to the best of our knowledge, there is no natural language processing technology that can process table information in 10-K documents, which means most numerical values of financial variables are completely ignored in textual analysis. So, textual information and financial variables complement each other very well.
@@
On the top of the dynamic industry network, we further introduced recurrent neural networks to model the inter-dependencies among M&A events. Specifically, for an acquirer, an event embedding $event_{d}^{t}\in\mathbb{R}^{3}$ is created as the vector representation of the information of the event happening in the timestamp $t$ in the focal acquirer $d$’s timeline. The four elements of an event embedding are: time difference $\Delta t$ between last event and the focal event, event type $I$, TNIC (Hoberg & Philips, 2016) 777see https://hobergphillips.tuck.dartmouth.edu/industryclass.htm for detail similarity $s$ between the acquirer and the target. There are two types of event that we consider in the event embeddings: self-triggered M&A event (represented by an integer 0) and peer effect (represented by an integer 1). For example, on the timeline of focal acquirer $d$, at timestamp $t$, a peer firm triggers a M&A event as acquirer. First, we look up what is the timestamp that the most recent event happened before this event, and compute the time difference $\Delta t$ between them. Then, since this is a M&A event triggered by a peer firm, the event type will be embedded as an integer value “1”. Finally, the TNIC similarity value is looked up in the TNIC dataset, and a corresponding event embedding is created. The event embedding will then be passed into a recurrent neural network as input.
@@
Given the big picture of the timing module as figure 7, if we view it in order of time, from past to recent, the multiple lines are a multi-dimensional point process while each line models the M&A behaviors of each acquirer. If we view it cross-sectionally, each slice is an industry network that directs the path of the propagation of peer effects. Different slices show different graph structures since the industry network is a dynamic network: the relationship between firms varies over time, which changes the direction of the propagation. In addition, for each acquirer, the RNN888the RNN shares parameters over all acquirers helps remember the peer effect from the past.
@@ 3.4 Data
For the M&A events data, we first remove all the M&A deals that happened within a company. They are asset reassignment and rearrangement inside a single firm which we do not consider. We further solely focus on complete and majority takeovers following the standard definition of a “merger and acquisition” (Betton et al., 2008; Ahern et al., 2014; Routledge et al., 2017). Specifically, the definition of “majority” takeover is the following: the percentage of share that acquirer acquire has to exceed 20% of the total share of the target; the fraction of shares held by the bidder after a completed transaction must be larger than 50%; the value of the transaction has to surpass a million US dollar; The transaction type could not be “block purchases”, “creeping acquisition”, and “privatization”. Similar to recommendation systems, we are facing cold-start problem (Schein et al., 2002; Lika et al., 2014) for those acquirers who never trigger any M&A events or only trigger a few. An analogy in recommendation systems in social networks is: if a user never links to any other user, it is nearly impossible to make precise recommendations to this user. Similarly, if a firm has never acquired any firm before, it is very hard to predict which firm will be targeted by it. To alleviate this issue, we only consider the firms who already triggered 4 or more merger and acquisition events during the observation window. We call them “frequent acquirers”, and we have 385 of them. For accounting variables, we majorly follow the choices of two recent M&A predictive works (Yang et al., 2014; Bernard et al., 2020), but also consider the data availability. Our final choices comprehensively cover the basic operation information for companies. They are variables about the assets of a firm, leverage, current condition of business operation, and the research and development spend131313see Appendix B for detail. Unfortunately, the missing ratios of the selected variables are still not low. For firms who occasionally have several missing values, we interpolate that value from the previous year. If a variable is totally missing over the entire observation period for a firm, we interpolate the average number of all other firms. For TNIC data that we used to construct the pre-defined edges in the industry network, we obtain the TNIC data from 1996 to 2019 (one year before the observation window). Since TNIC is conceptually a fully connected network with different weights on each edge, we only keep the top ten peers which follows the previous works (Hoberg & Philips, 2010; Hoberg & Philips, 2016). However, some industries are highly competitive, and others may not. To allow our network structures to have adaptability for different industries, we also set an absolute threshold for the TNIC similarity. In other words, if TNIC similarity exceeds a certain threshold, we will keep the link in the network while ignoring the former rule.
@@ 3.6 Evaluation
Proposed Model 0.5819 6.6%
@@ 4 Conclusion
The choice of accounting variables mainly follows two recent M&A predictive studies (Yang et al., 2014; Bernard et al., 2020). Yang et al. (2014) mentioned many financial, managerial and other variables141414see table 2 of (Yang et al., 2014) for more detail. For financial variables, they are: size, market-to-book ratio, cash, return on assets(ROA), sales to total assets (assets turnover), debt-to-equity ratio (leverage), current ratio, return on equity(ROE), debt-to-assets, capital expenditures to total asset, capital expenditures to total asset growth, sale growth, cash flow to total asset ratio, cash flow to sale ratio, asset growth. For managerial variables, they are: management inefficiency, resource richness, industry variations, relevance degree of business boundaries, export orientation, age of the firm. For other variables, they are: cost-to-income ratio, cost of goods sold divided by inventory, outperformance, tax shield effects, ratio of tangible assets to total assets, return on investment, profit margin, average dividend for last three years, shareholders, earnings before interest and taxes, operating income after depreciation, common shares traded divided by common shares outstanding. Bernard et al. (2020) mentioned seven accounting variables, the only one that was not mentioned by Yang et al. (2014) is PPE151515see definition at Appendix A of (Bernard et al., 2020) . However, due to the restriction of data availability, we only adopt part of the variables mentioned above. The variable we adopted, the formula for creating those variables, and the brief definitions of those variables are listed below [^16] The abbreviations follow Compustat..

### v3

# A Deep Learning Method for Predicting Mergers and Acquisitions: Temporal Dynamic Industry Networks

@@ Abstract
Merger and Acquisition (M&A) activities play an increasingly crucial role in market consolidation and restructuring. For companies pursuing acquisitions, M&A represents a significant investment strategy, with one of the primary incentives being the attainment of complementarities to enhance market power within competitive industries. Beyond the intrinsic factors extensively studied in existing empirical research, a firm’s M&A behavior is actively influenced by the M&A activities of its peer firms—a phenomenon known as the “peer effect.” However, current research often fails to capture the rich interdependencies of M&A events within the industry network.
An effective M&A predictive model should provide fine-grained, deal-level predictions without requiring ad-hoc feature engineering or data rebalancing. Such a model can deliver precise predictions of rival firms’ M&A behaviors and offer specific M&A recommendations for bidder and target companies. Most existing models, however, predict the behavior of only one side of M&A deals, lacking the capacity to make firm-specific deal recommendations. Additionally, they often rely on ad-hoc data truncations that convert continuous time information into arbitrary intervals, leading to information loss and impairing predictive performance. The laborious and arbitrary nature of these transformations also hampers the applicability of these models in real-world scenarios. Moreover, due to the sparsity of M&A events, existing deal-level predictive models often require ad-hoc data rebalancing, which introduces bias and further compromises performance and reliability.
To address these limitations, we propose a novel M&A predictive model based on Temporal Dynamic Industry Networks (TDIN), leveraging temporal point processes and deep learning techniques. Our approach captures the rich interdependencies among M&A events without the need for ad-hoc feature engineering or data rebalancing, as the temporal point processes inherently model event sparsity within their intensity functions. Empirical evaluations on M&A cases from January 1997 to December 2020 demonstrate the superiority of our method in predicting M&A activities and providing actionable deal-level recommendations. [^1] Code is available at: https://github.com/dayuyang1999/Merger_Acquisition_Prediction
@@ 2.2 Temporal Point Processes
We have the relationship between the intensity function and probability density function444proof see Appendix:
@@ 2.3 Graph Neural Network
The receptive field can be increased by stacking aggregation and combination mechanisms. If we have 2 layers of them, for every node, it can aggregate the information over 2-hop neighbors. However, generally speaking, the number of layers cannot be larger than 2 555see section 4.4 of (Hamilton et al., 2017) for detail. Aggregate and combine components are also called ”message passing mechanism”. Formally, consider the graph $G=(V,E)$, where $V$ is the set of nodes. The feature matrix of nodes is $X$. The aggregate and combine components at layer $l$ can be represented by:
@@ 3.2.1 Timing Module
Specifically, for the vector representation of intrinsic factors $e_{d}^{t}$, we consider two parts of the information as the input: accounting variables and textual embeddings from 10-K666only the “product description” part of 10-K. The reasons why we consider textual information are: first, recent studies have shown that textual information from 10-K has significant predictive power for M&A events (Hoberg & Philips, 2010; Hoberg & Philips, 2016; Routledge et al., 2017). Secondly, one of the most important motivations of firms continuing to conduct M&A transactions is that they are seeking product synergy to maintain competitive advantage (Hoberg & Philips, 2010). However, accounting variables have no capability to provide adequate information about firms’ products. In the meanwhile, to the best of our knowledge, there is no natural language processing technology that can process table information in 10-K documents, which means most numerical values of financial variables are completely ignored in textual analysis. So, textual information and financial variables complement each other very well.
@@
On the top of the dynamic industry network, we further introduced recurrent neural networks to model the inter-dependencies among M&A events. Specifically, for an acquirer, an event embedding $event_{d}^{t}\in\mathbb{R}^{3}$ is created as the vector representation of the information of the event happening in the timestamp $t$ in the focal acquirer $d$’s timeline. The four elements of an event embedding are: time difference $\Delta t$ between last event and the focal event, event type $I$, TNIC (Hoberg & Philips, 2016) 888see https://hobergphillips.tuck.dartmouth.edu/industryclass.htm for detail similarity $s$ between the acquirer and the target. There are two types of event that we consider in the event embeddings: self-triggered M&A event (represented by an integer 0) and peer effect (represented by an integer 1). For example, on the timeline of focal acquirer $d$, at timestamp $t$, a peer firm triggers a M&A event as acquirer. First, we look up what is the timestamp that the most recent event happened before this event, and compute the time difference $\Delta t$ between them. Then, since this is a M&A event triggered by a peer firm, the event type will be embedded as an integer value “1”. Finally, the TNIC similarity value is looked up in the TNIC dataset, and a corresponding event embedding is created. The event embedding will then be passed into a recurrent neural network as input.
@@
Given the big picture of the timing module as figure 7, if we view it in order of time, from past to recent, the multiple lines are a multi-dimensional point process while each line models the M&A behaviors of each acquirer. If we view it cross-sectionally, each slice is an industry network that directs the path of the propagation of peer effects. Different slices show different graph structures since the industry network is a dynamic network: the relationship between firms varies over time, which changes the direction of the propagation. In addition, for each acquirer, the RNN999the RNN shares parameters over all acquirers helps remember the peer effect from the past.
@@ 3.4 Data
For the M&A events data, we first remove all the M&A deals that happened within a company. They are asset reassignment and rearrangement inside a single firm which we do not consider. We further solely focus on complete and majority takeovers following the standard definition of a “merger and acquisition” (Betton et al., 2008; Ahern et al., 2014; Routledge et al., 2017). Specifically, the definition of “majority” takeover is the following: the percentage of share that acquirer acquire has to exceed 20% of the total share of the target; the fraction of shares held by the bidder after a completed transaction must be larger than 50%; the value of the transaction has to surpass a million US dollar; The transaction type could not be “block purchases”, “creeping acquisition”, and “privatization”. Similar to recommendation systems, we are facing cold-start problem (Schein et al., 2002; Lika et al., 2014) for those acquirers who never trigger any M&A events or only trigger a few. An analogy in recommendation systems in social networks is: if a user never links to any other user, it is nearly impossible to make precise recommendations to this user. Similarly, if a firm has never acquired any firm before, it is very hard to predict which firm will be targeted by it. To alleviate this issue, we only consider the firms who already triggered 4 or more merger and acquisition events during the observation window. We call them “frequent acquirers”, and we have 385 of them. For accounting variables, we majorly follow the choices of two recent M&A predictive works (Yang et al., 2014; Bernard et al., 2020), but also consider the data availability. Our final choices comprehensively cover the basic operation information for companies. They are variables about the assets of a firm, leverage, current condition of business operation, and the research and development spend141414see Appendix B for detail. Unfortunately, the missing ratios of the selected variables are still not low. For firms who occasionally have several missing values, we interpolate that value from the previous year. If a variable is totally missing over the entire observation period for a firm, we interpolate the average number of all other firms. For TNIC data that we used to construct the pre-defined edges in the industry network, we obtain the TNIC data from 1996 to 2019 (one year before the observation window). Since TNIC is conceptually a fully connected network with different weights on each edge, we only keep the top ten peers which follows the previous works (Hoberg & Philips, 2010; Hoberg & Philips, 2016). However, some industries are highly competitive, and others may not. To allow our network structures to have adaptability for different industries, we also set an absolute threshold for the TNIC similarity. In other words, if TNIC similarity exceeds a certain threshold, we will keep the link in the network while ignoring the former rule.
@@ 3.6 Evaluation
Proposed TDIN Model 0.5819 6.6%
4 Discussion and Analysis
To comprehensively evaluate the effectiveness of our proposed Temporal Dynamic Industry Network (TDIN) model for M&A prediction, we conducted an ablation study by systematically removing or modifying key components of the model. This section combines the ablation study results with an in-depth discussion, providing insights into the contributions of each component and their implications for M&A prediction.
4.1 Ablation Study Results
Table 2 summarizes the performance impact of each component in our model. We report the AUC scores and the percentage change in performance when each component is ablated.
Table 2: Ablation Study Results
Model Variant AUC Score Performance Change
Full TDIN Model 0.5819 -
Component Ablations:
   Without Textual Embeddings 0.5603 -3.71%
   Without Dynamic Industry Network 0.5492 -5.62%
   Without Peer Effect Modeling 0.5557 -4.50%
   Without RNN in Timing Module 0.5685 -2.30%
   Without GNN in Choice Module 0.5541 -4.78%
4.2 Analysis of Model Components
4.2.1 Impact of Textual Embeddings
Removing the textual embeddings derived from 10-K filings resulted in a 3.71% decrease in AUC score. This confirms the importance of incorporating unstructured textual information to capture qualitative factors such as product synergies, strategic intentions, and managerial outlooks that are not fully reflected in financial variables. Textual analysis allows the model to understand nuanced information, which is critical in M&A decisions where qualitative assessments often play a significant role.
4.2.2 Importance of Dynamic Industry Network
The dynamic industry network, modeling the evolving relationships between firms, proved to be the most critical component, with its removal causing a significant 5.62% drop in performance. This highlights the necessity of capturing the temporal dynamics of inter-firm relationships and the competitive landscape. By utilizing a dynamic network, the model can account for shifts in industry structure, entry and exit of firms, and changing competitive pressures, which are vital factors influencing M&A activity.
4.2.3 Contribution of Peer Effect Modeling
Excluding the peer effect component led to a 4.50% decrease in performance, emphasizing the influence of recent M&A events by peer firms on a focal firm’s M&A decisions. This finding aligns with empirical studies indicating that firms are more likely to engage in M&A activities following similar actions by competitors due to factors like competitive rivalry, industry trends, and herding behavior. Modeling peer effects allows the TDIN to capture these contagion effects, improving predictive accuracy.
4.2.4 Role of RNN in Timing Module
The recurrent neural network (RNN) in the timing module contributed to a 2.30% improvement in performance. While its impact is relatively modest compared to other components, the RNN effectively captures temporal dependencies and patterns in the timing of M&A events. This temporal modeling is essential for understanding not just whether a firm will engage in M&A, but also when it is likely to occur, which is crucial for dynamic prediction tasks.
4.2.5 Significance of GNN in Choice Module
The graph neural network (GNN) in the choice module is a key component, with its removal resulting in a 4.78% decrease in performance. The GNN enables the model to perform message passing and aggregation over the dynamic industry network, allowing firms to collect and react to information about their rivals. This mechanism effectively captures the interdependencies and strategic interactions among firms, which are fundamental in M&A decision-making processes.
4.3 Overall Performance Improvement
The proposed TDIN model achieves a 6.6% improvement in AUC score compared to the traditional Acquisition Likelihood Model, demonstrating the effectiveness of integrating dynamic network structures, textual information, and advanced deep learning techniques in M&A prediction. The significant performance gain underscores the model’s ability to capture the complex, multifaceted nature of M&A activities.
4.4 Implications for M&A Prediction and Research
The findings from our ablation study have several implications:
  • Holistic Modeling Approach: The substantial performance drops observed when key components are removed indicate that M&A prediction benefits from a holistic modeling approach that integrates financial data, textual information, network dynamics, and temporal patterns.
  • Importance of Dynamic Networks: Modeling the industry as a dynamic network is crucial for capturing the evolving competitive landscape and strategic interactions among firms. This approach can be extended to other domains where inter-entity relationships play a significant role.
  • Advancement in Deep Learning Applications: The effective use of RNNs and GNNs highlights the potential of deep learning architectures in handling complex temporal and relational data in financial prediction tasks.
4.5 Limitations and Future Research
While the TDIN model shows promising results, there are limitations to consider:
  • Data Availability: The model relies on comprehensive financial and textual data, which may not be available for all firms, particularly smaller or private entities.
  • Model Complexity: The integration of multiple deep learning components increases computational complexity and may pose challenges for scalability and real-time prediction.
  • Interpretability: Deep learning models often lack transparency, which can be a concern in financial contexts where explainability is important for decision-makers.
Future research could address these limitations by exploring methods for simplifying the model without significant loss of accuracy, enhancing interpretability through techniques like attention mechanisms, and extending the model to incorporate additional data sources such as news articles or social media sentiment.
@@ 5 Conclusion
The choice of accounting variables mainly follows two recent M&A predictive studies (Yang et al., 2014; Bernard et al., 2020). Yang et al. (2014) mentioned many financial, managerial and other variables151515see table 2 of (Yang et al., 2014) for more detail. For financial variables, they are: size, market-to-book ratio, cash, return on assets(ROA), sales to total assets (assets turnover), debt-to-equity ratio (leverage), current ratio, return on equity(ROE), debt-to-assets, capital expenditures to total asset, capital expenditures to total asset growth, sale growth, cash flow to total asset ratio, cash flow to sale ratio, asset growth. For managerial variables, they are: management inefficiency, resource richness, industry variations, relevance degree of business boundaries, export orientation, age of the firm. For other variables, they are: cost-to-income ratio, cost of goods sold divided by inventory, outperformance, tax shield effects, ratio of tangible assets to total assets, return on investment, profit margin, average dividend for last three years, shareholders, earnings before interest and taxes, operating income after depreciation, common shares traded divided by common shares outstanding. Bernard et al. (2020) mentioned seven accounting variables, the only one that was not mentioned by Yang et al. (2014) is PPE161616see definition at Appendix A of (Bernard et al., 2020) . However, due to the restriction of data availability, we only adopt part of the variables mentioned above. The variable we adopted, the formula for creating those variables, and the brief definitions of those variables are listed below [^17] The abbreviations follow Compustat..